OpenStudy (anonymous):

The tickets for a high school basketball game are $2.75 for students and $4.00 for adults. At Tuesday night's game, 132 tickets were sold for a total of $428.00. How many student tickets were sold? Is it 80?

OpenStudy (johnweldon1993):

Well...2 equations s + a = 132 2.50s + 4.00a = 428.00 solve top equation for s s = 132 - a substitute that in for 's' in the second equation 2.50(132 - a) + 4.00a = 428.00 distribute 330 - 2.50a + 4.00a = 428.00 combine like terms 1.50a = 98 divide both sides by 1.50 a = 65.33 plug this back into that first equation for 'a' s + a = 132 s + 65.33 = 132 -65.33 -65.33 s = 66.66 tickets Sold for students *I know it isn't a whole number but it is the answer*

OpenStudy (anonymous):

The choices are 78 80 50 52

OpenStudy (johnweldon1993):

what!? umm let me try this 2.50(78) + 4.00(54) ...411.00 not right 2.50(80) + 4.00(52)...408.00 not right 2.50(50) + 4.00(82)....453.00 not right 2.50(52) + 4.00(80)...450.00 not right None of those are the correct answer

OpenStudy (anonymous):

Oh! sorry, I had 2.50 instead of 2.75!

OpenStudy (anonymous):

@johnweldon1993 ^^

OpenStudy (johnweldon1993):

Ahh okay...lol I was going to say REALLY!? haha..then yes 80 would be the correct answer

OpenStudy (anonymous):

Ok, thanks man.